Circular No. 7290

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

S DORADUS P. Massey, Kitt Peak National Observatory, reports a dramatic change

in the spectrum of S Dor, the archtype of luminous blue variables (LBVs): “Spectra (resolution 0.1 nm; range 410–470 nm) taken with the 1.5-m telescope at Cerro Tololo Interamerican Observatory (CTIO) on Oct. 21, 23, and 24 superficially resemble those of an extreme F-type supergiant, with the strong absorption-line ‘pseudo photospheric’ spectrum characteristic of an LBV at visual maximum. Hβ is in emission, and Hγ and Hδ are in absorption, superposed upon broad emission; otherwise the spectrum contains only absorption lines. The spectrum is very similar to that observed in η Car during an outburst in 1893 (Walborn and Liller 1977, Ap.J. 211, 181), and that currently seen in the Hubble-Sandage (1953, Ap.J. 118, 353) LBV star ‘B’ = V0330 in M33 (Massey et al. 1996, Ap.J. 469, 629, fig. 8). However, such a spectral appearance of S Dor has not been previously reported in past decades of study. Rather, the optical spectrum has always been described as being dominated by emission when the star has been in a low state, or by P-Cyg lines when the star has been in a high state (cf. Thackeray 1964, MNRAS 129, 169; Wolf and Kaufer 1997, in Luminous Blue Variables, ASP Conf. Ser. 120, 26). For instance, a spectrum (range 375–510 nm) obtained by me three years ago with the CTIO 1.5-m on 1996 Oct. 28 shows a rich assortment of emission lines comprised primarily of hydrogen and singly-ionized metal lines, with only the upper Balmer lines and Mg II 448.1-nm in absorption, similar to that observed in AF And (in M31), and the Hubble-Sandage variable star ‘C’ (V0268) in M33, which are usually considered examples of ‘quiescent’ LBVs. The strong evolution of the spectral appearance of S Dor corresponds to only a modest change in its photometry, with somewhat redder colors, and a small increase in visual luminosity. I obtained UBV CCD data nearly simultaneously with the spectroscopy using the CTIO 0.9-m (1996 Nov. 1) and Curtis Schmidt (1999 Oct. 25) telescopes; the star has changed from V = 9.43, B − V = +0.18, U − B = −0.85 to V = 9.16, B − V = +0.28, U − B = −0.59. S Dor clearly merits further monitoring during the current Magellanic Cloud observing season.”

GRB 990712 Corrigendum. On IAUC 7225, line 7, for (IAUC 7721) read (IAUC

7221)

1999 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green